OpenStudy (anonymous):

integrate cos (x/4) from x=-pi to pi

OpenStudy (experimentx):

\[ \int_{-\pi}^{\pi} \cos (x/4)dx = 0\] For all even function you have \[ \int_{-a}^{a}f(x) dx = 0\] give that there is no infinite discontinuity.